# ICD-LM: Configuring Vision-Language In-Context Demonstrations by Language Modeling

### v2

@@ Abstract
This paper studies how to configure powerful In-Context Demonstration (ICD) sequences for a Large Vision-Language Model (LVLM) to solve Vision-Language tasks through In-Context Learning (ICL). After observing that configuring an ICD sequence is a mirror process of composing a sentence, i.e., just as a sentence can be composed word by word via a Language Model, an ICD sequence can also be configured one by one. Consequently, we introduce an ICD Language Model (ICD-LM) specifically designed to generate effective ICD sequences. This involves creating a dataset of hand-crafted ICD sequences for various query samples and using it to train the ICD-LM. Our approach, diverging from traditional methods in NLP that select and order ICDs separately, enables to simultaneously learn how to select and order ICDs, enhancing the effect of the sequences. Moreover, during data construction, we use the LVLM intended for ICL implementation to validate the strength of each ICD sequence, resulting in a model-specific dataset and the ICD-LM trained by this dataset is also model-specific. We validate our methodology through experiments in Visual Question Answering and Image Captioning, confirming the viability of using a Language Model for ICD configuration. Our comprehensive ablation studies further explore the impact of various dataset construction and ICD-LM development settings on the outcomes. The code is given in https://github.com/ForJadeForest/ICD-LM.
@@ 1 Introduction
With the escalation in model size [4, 45, 7, 6, 32, 44], the capabilities of Large Language Models (LLMs) have been largely enhanced. One salient emerging ability is In-Context Learning (ICL) [38, 39, 8], which can make LLMs quickly adapt to novel tasks using a few examples that are termed In-Context Demonstrations (ICDs). Crucially, ICL does not need gradient updates to model parameters, thereby circumventing challenges tied to additional data collection and model fine-tuning. Inspired by such success in NLP, researchers in the Vision-Language (VL) domain also design Large Vision-Language Models (LVLMs) capable of ICL [47, 33, 1, 2], paving the way for novel research and applications. Then an elemental question arising from these trained LVLMs is how to achieve optimal ICL performance across various vision-language tasks.
@@
Figure 1: (a) The traditional ICD configuration methods separately select and order the ICDs, leading to sub-optimal ICL performance. (b) Our ICD-LM enables the step-by-step generation of ICD configurations and simultaneously considers the selection of ICDs and the ordering of ICD sequences.
Before delving into ICL performance optimization in VL, we first retrospect the studies on ICL in NLP. While well-trained LLMs are identified as good In-Context Learners [4, 21], it has been observed that diverse ICD configurations heavily affect the ICL performance. For example, using different examples from the supporting set as ICDs [18, 9, 23], or even changing the orders of these ICDs [20, 48, 11], may lead to significant performance fluctuations. Driven by these findings, researchers design diverse heuristic-based (e.g., embedding similarity-based) retrieval methods [18, 28, 30] to select representative examples as the ICDs. Then these ICDs are reordered based on some principles, e.g., the Minimal Description Length [40] or Global and Local Entropy [20], for further improving the results.
While employing these methods to configure the ICDs yields consistent improvements, they still suffer from a few limitations. First, the independent operation of retrieval and reordering processes potentially yields sub-optimal results. For example, using similarity-based retrieval methods followed by reordering of ICD configuration may lead to the sub-optimal ICL performance, as shown in Fig. 1 (a) (i), where colors closer to the query sample (grey) in the figure indicate higher similarity. Second, as shown in Fig. 1 (a) (ii), researchers have noted that different LLMs exhibit sensitivities to particular ICD configurations [20], suggesting that the ICD configuration should be model-specific. Yet, most retrieval or reordering strategies remain model-agnostic, potentially leading to sub-optimal results.
Then if we want to directly extend these NLP strategies to VL, the above-mentioned challenges will be further exacerbated by the multimodal combinatorial complexity of vision and language data. For instance, although retrieving similar text samples typically yields favorable results in NLP, simply retrieving similar single-modal data, such as images, may not guarantee the best performance in VL tasks [41]. However, attempting to match both visual and textual modalities during retrieval will further magnify the time complexity. Additionally, VL tasks have more diverse problem structures than NLP, necessitating distinct retrieval or reordering strategies tailored to each VL task, thereby elevating the design and labor overhead.
In this study, we address these issues by introducing a novel approach for configuring ICDs to solve diverse VL tasks. Before delving into our approach, we reconsider the ideal process of constructing an ICD sequence. Given the query sample, we hope to subsequently select a series of ICDs that can maximize the ICL performance. This iterative selection mirrors the construction of a sentence wherein the word is selected one by one to ensure fluency. Recognizing this analogy, we propose to treat the ICD configuration as a specialized Language Model (LM). The only difference is that the vocabulary of this ICD-LM is not conventional words, but are the examples from the supporting set that will be used as ICDs.
We show the ICD-LM pipeline in Fig. 1 (b). To deploy this ICD-LM, a primary challenge arises that we do not have a ”ground-truth” dataset that indicates which examples or their orders can form an optimal ICD sequence for a specific LVLM-based in-context learner. To address it, we construct a dataset containing the pairs of a query and the corresponding ICD sequence. During construction, we use this frozen specific LVLM to measure whether an ICD sequence is good for the query or not. In this way, the constructed dataset becomes model-specific. Then we use this dataset to train a Transformer-based ICD-LM where each token in the vocabulary represents an example from the supporting set. During training, the ICD-LM concurrently learns the selection and ordering of ICDs, eliminating the necessity to operate two separate stages for constructing an ICD sequence as previous methods.
Besides alleviating the above-mentioned limitations, we also find some other interesting characteristics of our method. First, we evaluate the length extrapolation ability of ICD-LM, which pertains to its performance when generating ICD sequences that exceed the maximum length represented in the dataset. This is contrasted with its interpolation ability,i.e., the performance within the maximum ICD length of that dataset. For instance, if a dataset only contains 2-shot ICDs, the interpolation ability denotes the effect of the 1- and 2-shot ICDs when implementing ICL, while the extrapolation ability denotes the effect of more shot ICDs like 4- or even 8-shot ICDs. Interestingly, we find that the trained ICD-LM has strong length extrapolation ability, e.g., when trained on a dataset containing only 2-shot ICDs, it can proficiently configure 4-shot ICDs. Second, the trained ICD-LM may configure a fixed ICD sequence for each query while achieving excellent ICL performance. For example, in the IC/VQA tasks, the golden ICD sequence achieves 6.91/1.24 improvements compared with a strong baseline that retrieves the ICDs by similar images.
@@ 2 Related Work
Models with In-Context Learning Ability. Prompt engineering enables pre-trained Large Language Models(LLMs) to address downstream tasks without the need for fine-tuning [24, 4, 29]. A variant, ICL, enhances this ability by constructing prompts with a few examples. This has been demonstrated in LLMs such as GPT-3 [4], LLaMA [32], and MPT [31]. Recently, witnessing such success in NLP, the VL domain has also developed numerous LVLMs with prompt engineering abilities [49, 17, 13, 1, 33, 43]. With the support of LLM, Frozen [33] and Flamingo [1] have developed the ability for ICL. Among these LVLMs, Flamingo simulates the ICL approach during training, resulting in a more powerful ICL ability. However, due to the unavailability of an official open-source version of Flamingo, we use an unofficial implementation, OpenFlamingo [2] in this study to explore whether ICD-LM can configure useful ICD sequences.
@@
Configuring In-Context Demonstrations. Although ICL assists LLMs in better adapting to downstream tasks, its performance is highly sensitive to the selection [18, 9, 23] and ordering [20, 48, 11] of ICDs. Numerous studies have explored diverse methods to select ICDs in the NLP field. For example,  [18] selects ICDs based on the embedding similarity between ICDs and test samples where the embeddings are extracted from an existent language encoder. Such method is further developed by training an encoder specifically for selection [26, 15, 42, 36, 15, 37, 5]. Additionally,  [40, 22, 14] leverage LLM outputs to select and  [46] formulates ICD selection as a Markov decision process [3].
Regarding the ordering of ICDs, researchers propose some principles to measure the quality of ICD configurations, e.g., the Minimal Description Length [40] and Global and Local Entropy [20]. However, these reordering methods are limited to classification tasks in NLP, as they require the calculation of conditional probabilities for each label given the task inputs. These approaches are inapplicable to certain tasks lacking a predefined label space, such as open-ended question answering or text generation.
@@
Besides these NLP studies, in VL, [41] explores diverse ICD configurations on image captioning, while only the heuristic-based methods are used for selecting ICDs and do not consider the ordering. In contrast, our ICD-LM can simultaneously learn how to select and reorder the samples and moreover, our ICD-LM is model-specific.
3 Model-Specific ICD-LM
@@
In this section, we introduce how to build a model-specific ICD-LM for configuring the ICD sequence for a given LVLM. First, we briefly introduce the formulations of ICL for Vision-Language (VL) tasks in section 3.1. Then we show how to construct a model-specific dataset (section 3.2) that will be used to train the ICD-LM, whose structure will be detailed in Section 3.3.
3.1 In-Context Learning (ICL)
@@
Given a query input $\mathbf{x}^{\prime}$, ICL predicts the corresponding output $\mathbf{y}^{\prime}$ using a well-trained foundation model $\mathcal{M}$, conditioned on the concatenation of an in-context sequence $\mathcal{S}$ and this query. If the in-context sequence contains $K$-shot ICDs, it is denoted it as: $\mathcal{S}^{K}=\{\hat{\bm{d}}_{1},\hat{\bm{d}}_{2},...,\hat{\bm{d}}_{K}\}$, where $\hat{\bm{d}}_{k}$ is the $k$-th ICD. Then ICL can be formulated as:
@@
where $\bm{P}_{\mathcal{M}}$ denotes the predicted probability of $\mathcal{M}$ and $\leftarrow$” represents the decoding strategy, e.g., beam search.
For each ICD $\hat{\bm{d}}$, it is selected from a supporting dataset $\mathcal{D}_{S}=\{\bm{d}_{1},...,\bm{d}_{N}\}$, where each sample $\bm{d}_{i}=(\mathbf{x}_{i},\mathbf{y}_{i})$: $\mathbf{x}_{i}$ and $\mathbf{y}_{i}$ respectively denote the input and the corresponding label. It is noteworthy that in diverse VL tasks, $\mathbf{x}$ and $\mathbf{y}$ have different forms. For instance, in Image Captioning (IC), $\mathbf{x}$ is the image and $\mathbf{y}$ is the caption; and in Vision Question Answering (VQA), $\mathbf{x}$ contains the image and the question, while $\mathbf{y}$ is the answer.
3.2 Constructing the Model-Specific Dataset
Figure 2: The pipeline of constructing $\mathcal{D}_{\mathcal{M}}$. The darker color of $S^{K}_{i,j}$ indicates a higher score given by Eq. 2 in figure (c).
To train our model-specific ICD-LM for generating effective ICD sequences for a given LVLM $\mathcal{M}$, we should first construct a model-specific dataset $\mathcal{D}_{\mathcal{M}}$ containing high-quality ICD sequences for different query inputs. To avoid confusion, Figure 2 shows how to construct $\mathcal{D}_{\mathcal{M}}$ from an VL dataset, e.g., the COCO dataset [16] for IC. Formally, given an dataset $\mathcal{D}$ which is already split into the training part $\mathcal{D}_{R}$ and the test part $\mathcal{D}_{E}$, we build $\mathcal{D}_{\mathcal{M}}$ only from the training part $\mathcal{D}_{R}$.
As shown in Fig. 2 (a), initially, we randomly select $n$ samples from $\mathcal{D}_{R}$ to form an anchor set $\mathcal{A}$. Then for each sample $\bm{a}_{m}=\{\bm{x}_{m},\bm{y}_{m}\}\in\mathcal{A}$, we aim to construct a $K$-shot in-context sequence $\mathcal{S}_{m}^{K}$ for it. Then $\mathcal{D}_{\mathcal{M}}=\{(\bm{a}_{1},\mathcal{S}_{1}^{K}),(\bm{a}_{2},\mathcal{S}_{2}^{K}),...,(\bm{a}_{M},\mathcal{S}_{M}^{K})\}$ where each training sample contains an query $\bm{a}_{m}$ and the corresponding $K$-shot in-context sequence $\mathcal{S}_{m}^{K}$.
To avoid confusion, we remove the subscript $m$ in following. To construct $\mathcal{S}^{K}=\{\bm{d}_{1},...,\bm{d}_{K}\}$, we need to select $K$-shot samples from the supporting set $\mathcal{D}_{S}$, which is set to the complement set of $\mathcal{A}$ in $\mathcal{D}_{R}$: $\mathcal{D}_{R}\backslash\mathcal{A}$. Meantime, we also need to decide which samples should be selected in turn. To achieve this, given the anchor sample $\bm{a}=\{\bm{x},\bm{y}\}$ and the partially constructed in-context sequence, e.g., a $k-1$-shot $\mathcal{S}^{k-1}$, we need to know that after adding which sample $\bm{d}\in\mathcal{D}_{S}$, the ICL performance improvement can be maximized by applying the given LVLM $\mathcal{M}$:
@@ 3.2 Constructing the Model-Specific Dataset
where $I_{\mathcal{M}}$ is one kind of ICL performance measurement related to $\mathcal{M}$. Note that Eq. (2) actually uses the greedy sampling method to select the samples every time, while we can use beam search here to further achieve a better solution. Additionally, to improve the diversity of the dataset, we will keep the top-$b$ highest-scoring ICD sequences $\{S^{K}_{m,1},S^{K}_{m,2},...,S^{K}_{m,b}\}$ at the last iteration for each $a_{m}$, where $b$ is equal to the beam size. For example, when setting beam size to $5$ as shown in Fig. 2 (c), we can get $5$ diverse high-quality ICD sequences for an anchor sample.
Intuitively, for diverse tasks, we can use the corresponding “golden measurement” as $I_{\mathcal{M}}$, e.g., setting it to CIDEr [35] for IC and accuracy for VQA. However, this strategy encounters two limitations. The first one is that for diverse VL tasks, we need diverse corresponding measurements, which is inconvenient. Second, some “golden measurements” may be impractical to deploy. For example, for IC, calculating CIDEr requires the LVLM to forward multiple times to sample an integral sentence, and then it costs expensive time burdens to construct the dataset. While for VQA, accuracy is a binary value (accuracy=1 when correct and 0 when wrong), then maybe lots of candidates in $\mathcal{D}_{S}$ will make the accuracy change from 0 to 1 and then it is hard to judge which one of them is the most suitable one.
@@
$I_{\mathcal{M}}(\mathcal{S}^{K},\bm{a})$ $=P_{\mathcal{M}}(\bm{y}|\mathcal{S}^{K},\bm{x})$ (3)
$=\prod_{t}P_{\mathcal{M}}(y^{(t)}|\mathcal{S}^{K},\bm{x},y^{(1:t-1)}).$
Note that in VL tasks, the ground-truth label $\bm{y}=\{y^{(1)},...,y^{(T)}\}$ is usually a text sequence, thus we can decompose the probability distribution into a series of productions. In this way, Eq. (2) selects a sample that can further maximize the prediction confidence given the query input and the current in-context sequence.
In implementation, $\mathcal{D}_{S}$ usually contains huge amounts of samples, e.g., $\mathcal{D}_{S}$ in COCO [16] contains about $10^{5}$ samples. However, we need to calculate Eq. (2) for each $\bm{d}\in\mathcal{D}_{S}$ when selecting $\hat{\bm{d}}_{k}$ for each $\bm{a}\in\mathcal{A}$, which means the whole process of building $\bm{D}_{\mathcal{M}}$ is quite time-consuming. To alleviate the cost, as shown in Fig. 2 (b), for each specific $\bm{a}$, we narrow the set size by sampling a much smaller subset $\mathcal{D}_{S}^{\bm{a}}$, e.g., containing 64 samples $\mathcal{D}_{S}^{\bm{a}}=\{d_{\bm{a}}^{1},d_{\bm{a}}^{2},...,d_{\bm{a}}^{64}\}$, from $\mathcal{D}_{S}$ for selecting $\hat{\bm{d}}_{k}$. We use diverse sampling strategies to construct this subset, e.g., retrieving some similar samples as $\bm{a}$, and implement exhaustive ablation studies to explore which strategies are useful in Section 4.4.
3.3 Training Model-Specific ICD-LM
Figure 3: Top: The architecture of ICD-LM, which is a two-layer Transformer decoder blocks. Bottom: The components of the input embeddings, which is the sum of the random initialized learnable embeddings, the image and text embeddings extracted by CLIP. The dotted block means that some tasks do not exist the text input, e.g., IC.
After constructing the model-specific dataset $\mathcal{D}_{\mathcal{M}}$, we use it to train our ICD-LM, as Figure 3 shows, it is a two-layer Transformer decoder blocks [34]. The primary difference between our ICD-LM and the traditional LM lies in the tokens of the vocabulary. Specifically, the tokens in our vocabulary are the samples from the supporting set $\mathcal{D}_{S}$, e.g., the first token corresponds to the first sample in $\mathcal{D}_{S}$.
@@ 3.3 Training Model-Specific ICD-LM
To train the ICD-LM, we first need to embed the tokens of the vocabulary to get dense embeddings. Since each token contains both image and text, we use the vision encoder $F_{I}(\cdot)$ and the language encoder $F_{T}(\cdot)$ of CLIP [25] to embed the image and text, respectively. Meanwhile, we add each of these embeddings with a learnable part which is randomly initialized. Then for the $i$-th token $\bm{d}_{i}=(I_{i},T_{i})$ in the vocabulary where $I_{i}$ and $T_{i}$ are the corresponding image and text, respectively, its token embedding is $\bm{e}_{i}$:
$\bm{e}_{i}=F_{I}(I_{i})+F_{T}(T_{i})+\bm{r}_{i},$ (4)
where $\bm{r}_{i}$ is the randomly initialized learnable embedding for the $i$-th token. Note that $T_{i}$ varies between IC and VQA tasks. In IC, $T_{i}$ represents the caption, whereas in VQA, it denotes the question and the answer.
For the test query $x^{\prime}$, we also use the same vision and language encoders to embed it. For VQA, the image and question will be embedded and summed, while for IC, only the image is embedded. Lastly, we employ the cross-entropy loss for training as a standard LM that given the previously $k-1$ tokens, we maximize the probability of the $k$-th ground-truth token.
3.4 Configuring the ICD Sequence

### v4

# Lever LM: Configuring In-Context Sequence to Lever Large Vision Language Models

@@ Abstract
As Archimedes famously said, “Give me a lever long enough and a fulcrum on which to place it, and I shall move the world”, in this study, we propose to use a tiny Language Model (LM), e.g., a Transformer with 67M parameters, to lever much larger Vision-Language Models (LVLMs) with 9B parameters. Specifically, we use this tiny Lever-LM to configure effective in-context demonstration (ICD) sequences to improve the In-Context Learinng (ICL) performance of LVLMs. Previous studies show that diverse ICD configurations like the selection and ordering of the demonstrations heavily affect the ICL performance, highlighting the significance of configuring effective ICD sequences. Motivated by this and by re-considering the the process of configuring ICD sequence, we find this is a mirror process of human sentence composition and further assume that effective ICD configurations may contain internal statistical patterns that can be captured by Lever-LM. Then a dataset with effective ICD sequences is constructed to train Lever-LM. After training, given novel queries, new ICD sequences are configured by the trained Lever-LM to solve vision-language tasks through ICL. Experiments show that these ICD sequences can improve the ICL performance of two LVLMs compared with some strong baselines in Visual Question Answering and Image Captioning, validating that Lever-LM can really capture the statistical patterns for levering LVLMs. The code is available at https://github.com/ForJadeForest/Lever-LM.
@@ 1 Introduction
With the escalation in model size and training data [1, 2, 3, 4, 5, 6], Large Language Models (LLMs) emerge the ability of In-Context Learning (ICL) [7, 8, 9]. ICL, akin to few-shot learning [10, 11, 12], utilizes a few exemplary In-Context Demonstrations (ICDs) to adapt LLMs to new tasks without gradient updates. This achievement in NLP has inspired researchers to similarly enhance Large Vision-Language Models (LVLMs) with ICL capabilities [13, 14]. However, just as in NLP, the effectiveness of ICL in LVLMs is significantly influenced by the configurations of ICDs, such as their selection and ordering [15, 16, 17, 18, 19, 20, 21, 22, 23]. Recent studies [24, 25, 26] have shown that this sensitivity in LVLMs is further exacerbated by the multimodal combinatorial complexity of vision and language data.
In NLP, researchers employ various strategies to optimize in-context sequences to improve ICL performance, including retrieving representative examples as the ICDs [15, 27, 28] and re-ordering these ICDs based on specific principles [29, 21]. While these methods have shown improvements, their application remains largely confined to NLP and is less explored in the vision-language domain. Moreover, as shown in Fig. 1(a), the independent operations of retrieval and reordering often result in sub-optimal outcomes. A critical reconsideration of the ICD sequence generation reveals that configuring an optimal ICD sequence should be a coherent process. Instead of independently selecting and re-ordering, each ICD should be chosen conditionally based on the previous ICDs. This mirrors the sequential nature of human sentence composition, where each word is sequentially selected to ensure overall fluency. Such fluency can be characterized by temporal statistical patterns, which allows the design of statistical learning methods to model and learn from data, where Language Model is one typical technique that demonstrates the effectiveness. This analogy supports the hypothesis that optimal ICD sequences may contain inherent temporal statistical patterns.
@@
Figure 1: (a) The traditional ICD configuration methods separately select and order the ICDs, leading to sub-optimal ICL performance. (b) Our Lever-LM enables the step-by-step generation of ICD configurations and simultaneously considers the selection of ICDs and the ordering of ICD sequences.
Although not explicitly stated, some previous studies in NLP work toward this direction by calculating statistical metrics like the perplexity [30] or the entropy [31, 32] to discover what statistical characteristics a good prompt should have. However, the requirement to know the probability of each token when calculating these metrics limits their applicability in VL. This is because recently proposed LVLMs [13, 14, 33] use continuous image patches rather than the tokenized discrete elements as the vision input, rendering these techniques used in NLP inapplicable in VL and the following two questions remain unaddressed: (1) whether effective VL ICDs exhibit certain statistical patterns and (2) whether such patterns can be leveraged to compose new ICD sequences for a given query. This study aims to address these questions. Specifically, we employ a tiny Language Model, e.g., a Transformer, to capture the inherent statistic patterns. This tiny LM is named as “Lever-LM” since it can lever/control a much larger VLM by composing suitable ICD sequence. Compared with the classic LM, the only difference is that the vocabulary of Lever-LM consists not of standard words, but of examples from the supporting set that will be used as ICDs.
Fig. 1 (b) shows the training pipeline for Lever-LM. Initially, a "ground-truth" dataset is constructed to indicate which examples or their orders can form good ICD sequences. Specifically, we employ a frozen LVLM to evaluate if an ICD sequence facilitates accurate predictions for a given query, e.g., answering questions correctly or generating appropriate captions.  [^1] Note: This method requires ground-truth labels and thus can not be used at the test stage. Lever-LM is then trained to concurrently learn the selection and ordering of ICDs, streamlining the process by eliminating the need for two separate stages typical of previous methods. Our experiments, conducted with two LVLMs—Open-Flamingo [34] and IDEFICS [14]—on classic VL tasks—Image Captioning and Visual Question Answering— demonstrate that Lever-LM surpasses several strong baselines, including those that retrieve ICDs based on image similarity. These results confirm that effective ICD sequences contain inherent temporal statistical patterns and such patterns can be learned for composing new ICD sequences for test queries.
Besides the above-mentioned advantages, Lever LM emerges two interesting abilities. First, it has strong length extrapolation ability, e.g., when trained on a dataset with only 2-shot ICDs, Lever LM can generate 4 or more-shot ICDs that outperform several strong baselines. Second, Lever LM can construct a “golden” ICD sequence of 8 predetermined ICDs in a fixed order. This sequence can be uniformly applied across different test queries to assist LVLM in label generation, thereby reducing the computational overhead for configuring new ICD sequences for each query. Experiments in IC/VQA tasks show that “golden” ICD sequence achieves 6.91/1.24 improvements compared to a strong baseline. In addition, we use exhaustive ablations, including applying different ways to construct training set and changing the architecture of Lever-LM, to discover which factors and analyze why they will affect the ICL performance.
@@ 2 Related Work
Models with In-Context Learning Ability. Prompt engineering enables Large Language Models (LLMs) to address downstream tasks without the need for fine-tuning [35, 1, 36]. A variant, ICL, enhances this ability by constructing prompts with a few examples. This has been demonstrated in LLMs such as GPT-3 [1], LLaMA [5], and MPT [37]. Recently, witnessing such success in NLP, the VL domain has also developed numerous LVLMs with prompt engineering abilities [38, 33, 39, 13, 40, 41, 42, 43, 44] and ICL ability like  [40], Flamingo [13], and IDEFICS [14] Among them, we use Flamingo and IDEFICS as the LVLMs to explore the effectiveness of Lever-LM since they have stronger and more robust ICL ability by using better language encoders and more training data222 Since Flamingo does not open-source the model, we use an unofficial implementation, OpenFlamingo [34]..
@@
Configuring In-Context Demonstrations. Although ICL assists LLMs in better adapting to downstream tasks, its performance is highly sensitive to the selection [15, 19, 20] and ordering [21, 22, 23] of ICDs. Numerous studies have explored diverse methods to select ICDs in the NLP field  [29, 45, 46, 47, 48]. For example,  [15] selects ICDs based on the embedding similarity between ICDs and test samples where the embeddings are extracted from an existing language encoder. Such a method is further developed by training an encoder specifically for selection [49, 50, 51, 52, 50, 53].
Regarding the ordering of ICDs, researchers calculate diverse statistical-based metrics to measure the quality of ICD configurations, e.g., the Minimal Description Length [29] and Global and Local Entropy [21]. Besides them, researchers focus more on discovering the statistical patterns of good prompts. For example, [30] uses perplexity to measure which prompts can better help LLMs perform a task. Furthermore,  [31] unifies diverse statistics-based prompt selection methods [54, 55] from the perspective of mutual information and discover that mutual information or its variants can uncover certain statistical patterns of effective prompts. However, these statistical-based methods require to calculate the token probabilities, making them infeasible to address continuous image patches, thus can not be used in VL.
@@
Besides these NLP studies, in VL, [24] and [25] explore diverse ICD configurations in IC and VQA, while only the heuristic-based methods are used for selecting ICDs and do not consider the ordering. In contrast, our Lever-LM can simultaneously learn how to select and reorder the samples and moreover, our Lever-LM is model-specific.
3 Lever Language Model
@@
In this section, we introduce how to build Lever Language Model (Lever-LM) for configuring the ICD sequence to lever a given LVLM. First, we briefly introduce the formulations of ICL for Vision-Language (VL) tasks. Then we introduce the construction of the dataset used to train Lever-LM. After that, we show the architecture of Lever-LM and briefly discuss how to train Lever-LM and use it to configure ICD sequences.
@@
The Formulation of In-Context Learning (ICL). Given a query input $\mathbf{x}^{\prime}$, ICL predicts the corresponding output $\mathbf{y}^{\prime}$ using a well-trained foundation model $\mathcal{M}$, conditioned on the concatenation of an in-context sequence $\mathcal{S}$ and this query. We denote a in-context sequence with $K$-shot ICDs $\hat{\bm{d}}$ as $\mathcal{S}^{K}=\{\hat{\bm{d}}_{1},\hat{\bm{d}}_{2},...,\hat{\bm{d}}_{K}\}$. Then ICL can be formulated as:
@@
where $\bm{P}_{\mathcal{M}}$ denotes the predicted probability of $\mathcal{M}$ and $\leftarrow$” represents the decoding strategy, e.g., beam search. For each $\hat{\bm{d}}$, it is selected from a supporting set $\mathcal{D}_{S}=\{\bm{d}_{1},...,\bm{d}_{N}\}$, where each sample $\bm{d}_{i}=(\mathbf{x}_{i},\mathbf{y}_{i})$: $\mathbf{x}_{i}$ and $\mathbf{y}_{i}$ respectively denote the input and the corresponding label. It is noteworthy that in diverse VL tasks, $\mathbf{x}$ and $\mathbf{y}$ have different forms. For instance, in Image Captioning (IC), $\mathbf{x}$ is the image and $\mathbf{y}$ is the caption; and in Vision Question Answering (VQA), $\mathbf{x}$ contains the image and the question, while $\mathbf{y}$ is the answer.
Figure 2: (a): The pipeline of constructing $\mathcal{D}_{\mathcal{M}}$. Darker color of $S^{K}_{i,j}$ indicates a higher score given by Eq. 2. (b): Top: Lever-LM is a two-layer Transformer. Bottom: Each input embeddings is the sum of the random initialized learnable embeddings, the image and text embeddings extracted by CLIP. The dotted block means that some tasks do not exist the text input, e.g., IC.
Constructing the Training Dataset. To train Lever-LM for generating effective ICD sequences for a given LVLM $\mathcal{M}$, we should first construct a dataset $\mathcal{D}_{\mathcal{M}}$ containing high-quality ICD sequences for different query inputs. Simply, we use one VL dataset—COCO [56] from IC— to show how to construct $\mathcal{D}_{\mathcal{M}}$, which is shown in Fig. 2. Formally, given a dataset $\mathcal{D}$ which is already split into the training part $\mathcal{D}_{R}$ and the test part $\mathcal{D}_{E}$, we build $\mathcal{D}_{\mathcal{M}}$ only from $\mathcal{D}_{R}$. As Fig. 2 (a.1) shows, initially, we randomly select $n$ samples from $\mathcal{D}_{R}$ to form an anchor set $\mathcal{A}$. Then for each sample $\bm{a}_{m}=\{\bm{x}_{m},\bm{y}_{m}\}\in\mathcal{A}$, we construct a $K$-shot in-context sequence $\mathcal{S}_{m}^{K}$ for it. Then $\mathcal{D}_{\mathcal{M}}=\{(\bm{a}_{1},\mathcal{S}_{1}^{K}),(\bm{a}_{2},\mathcal{S}_{2}^{K}),...,(\bm{a}_{M},\mathcal{S}_{M}^{K})\}$ where each training sample contains an query $\bm{a}_{m}$ and the corresponding $K$-shot in-context sequence $\mathcal{S}_{m}^{K}$.
To avoid confusion, we remove the subscript $m$ in following texts. To construct $\mathcal{S}^{K}=\{\bm{d}_{1},...,\bm{d}_{K}\}$, we need to select $K$-shot samples from the supporting set $\mathcal{D}_{S}$, which is set to the complement set of $\mathcal{A}$ in $\mathcal{D}_{R}$: $\mathcal{D}_{R}\backslash\mathcal{A}$. Meantime, we also need to decide which samples should be selected in turn. To achieve this, given the anchor sample $\bm{a}=\{\bm{x},\bm{y}\}$ and the partially constructed in-context sequence, e.g., a $k-1$-shot $\mathcal{S}^{k-1}$, we need to know that after adding which sample $\bm{d}\in\mathcal{D}_{S}$, the ICL performance improvement can be maximized by applying the given LVLM $\mathcal{M}$:
@@
where $I_{\mathcal{M}}$ is one kind of ICL performance measurement related to $\mathcal{M}$. Note that Eq. (2) actually uses the greedy sampling method to select the samples every time, while we can use beam search here to further achieve a better solution. Additionally, to improve the diversity of the dataset, we will keep the top-$b$ highest-scoring ICD sequences $\{S^{K}_{1},S^{K}_{2},...,S^{K}_{b}\}$ at the last iteration for each $a$, where $b$ is equal to the beam size. For example, when setting beam size to $5$ as shown in Fig. 2 (a.3), we can get $5$ diverse high-quality ICD sequences for an anchor sample.
Intuitively, for diverse tasks, we can use the corresponding “golden measurement” as $I_{\mathcal{M}}$, e.g., setting it to CIDEr [57]/accuracy for IC/VQA. However, this strategy encounters two limitations. The first one is that for diverse VL tasks, we need diverse corresponding measurements, which is inconvenient. Second, some “golden measurements” may be impractical to deploy. For example, for IC, calculating CIDEr requires the LVLM to forward multiple times to sample an integral sentence, and then it costs expensive time burdens to construct the dataset. While for VQA, accuracy is a binary value (accuracy=1 when correct and 0 when wrong), then maybe lots of candidates in $\mathcal{D}_{S}$ will make the accuracy change from 0 to 1 and then it is hard to judge which one of them is the most suitable one.
@@
$\small I_{\mathcal{M}}(\mathcal{S}^{K},\bm{a})=P_{\mathcal{M}}(\bm{y}|\mathcal{S}^{K},\bm{x})=\prod\nolimits_{t}P_{\mathcal{M}}(y^{(t)}|\mathcal{S}^{K},\bm{x},y^{(1:t-1)}).$ (3)
In VL tasks, the ground-truth label $\bm{y}=\{y^{(1)},...,y^{(T)}\}$ is a sequence, thus we can decompose the probability distribution into a series of productions. Then Eq. (2) selects a sample that can further maximize the prediction confidence given the query input and the current in-context sequence.
In implementation, $\mathcal{D}_{S}$ usually contains huge amounts of samples, e.g., $\mathcal{D}_{S}$ in COCO [56] contains about $10^{5}$ samples. However, we need to calculate Eq. (2) for each $\bm{d}\in\mathcal{D}_{S}$ when selecting $\hat{\bm{d}}_{k}$ for each $\bm{a}\in\mathcal{A}$, which means the whole process of building $\bm{D}_{\mathcal{M}}$ is quite time-consuming. To alleviate the cost, as shown in Fig. 2 (a.2), for each specific $\bm{a}$, we narrow the set size by sampling a much smaller subset $\mathcal{D}_{S}^{\bm{a}}$, e.g., containing 64 samples $\mathcal{D}_{S}^{\bm{a}}=\{d_{\bm{a}}^{1},d_{\bm{a}}^{2},...,d_{\bm{a}}^{64}\}$, from $\mathcal{D}_{S}$ for selecting $\hat{\bm{d}}_{k}$. We use diverse sampling strategies to construct this subset, e.g., retrieving some samples similar to $\bm{a}$, and implement exhaustive ablation studies to explore which strategies are useful in Section 4.3.
Training Lever-LM. After getting $\mathcal{D}_{\mathcal{M}}$, we use it to train Lever-LM, as Fig. 2(b) shows, it is a two-layer tiny Transformer [58]. The primary difference between Lever-LM and the traditional LM lies in the tokens of the vocabulary, whose tokens are the samples from the supporting set $\mathcal{D}_{S}$, e.g., the first token corresponds to the first sample in $\mathcal{D}_{S}$. Then given the query sample, the ICDs can be selected one by one based on the token distribution produced by the trained Lever-LM, just as when composing a sentence, the words are selected one by one from the word vocabulary.
@@
To train Lever-LM, we should embed the tokens of the vocabulary to get dense embeddings. Since each token contains both image and text, we use the vision encoder $F_{I}(\cdot)$ and the language encoder $F_{T}(\cdot)$ of CLIP [59] to embed the image and text, respectively. Meanwhile, we add each of these embeddings with a learnable part $\bm{r}_{i}$ that is randomly initialized. Then for the $i$-th token $\bm{d}_{i}=(I_{i},T_{i})$ in the vocabulary where $I_{i}$/$T_{i}$ are the corresponding image/ text, its token embedding is $\bm{e}_{i}$:
$\bm{e}_{i}=F_{I}(I_{i})+F_{T}(T_{i})+\bm{r}_{i}.$ (4)
Note that $T_{i}$ varies between IC and VQA tasks where it denotes caption in IC and question in VQA.
For test query $x^{\prime}$, we use the same vision and language encoders to embed it. For VQA, the image and question are embedded and summed, while for IC, only the image is embedded. Lastly, we use the cross-entropy loss for training as a standard LM that given the previously $k-1$ tokens, we maximize the probability of the $k$-th ground-truth token.